# Cyber Food Swamps: Investigating the Impacts of Online-to-Offline Food Delivery Platforms on Healthy Food Choices

### v2

@@ Introduction
By comparing the healthiness of online and offline food environments with fast food preferences in regional O2O food delivery orders, we find that the correlation is stronger for online environments ($r=0.204$) compared to offline environments ($r=0.092$). Figure 7 demonstrates that variations in online food environments are associated with more distinct differences in healthy food choices. These findings suggest notable differences between online and traditional offline food environments. The stronger relationship between the online food environment and healthy food choices underscores the need for further investigation into their causal connections.
@@
For the online food order decision $y_{it}$ by user $i$ located in region $c_{it}$ at time $t$, we characterize the online food environment using $\phi(c_{it})$ (the proportion of accessible fast food restaurants), and user $i$’s fast food preference $\psi_{i}(t)$, as the fraction of fast food orders made by user $i$ in the six months prior to time $t$. We focus on 91,879 users who have made at least 20 orders, extracting 3.86 million orders to fit a logistic regression model to estimate the effect of online food environment $\phi(c_{it})$ on the decision to choose fast food over non-fast food:
@@
where $y_{it}=1$ represents ordering fast food, $\text{logit}^{-1}(x)=\frac{e^{x}}{1+e^{x}}$ is the logistic function, $\beta_{0}$ is the fixed intercept term, and $\delta_{t}$ is a fixed effect term accounting for monthly variation. $\beta_{i}$ and $\beta_{c}$ characterize the separate effect of user preference and online food environment, respectively, as shown in Figure 8.
Figure 8: Effect of user’s historical preference ($\beta_{i}$) and online food environment ($\beta_{c}$) on the probability of placing a fast food delivery order over a non-fast food delivery order for different gender, income, and age groups. Error bars represent 95% confidence intervals for $\beta$’s.
The regression model reveals that both online food environment and personal preference are significantly associated with fast food choice, with personal preference exhibiting a higher log-odd coefficient ($\beta_{i}=3.227\pm 0.012$, $P<0.001$) than the online food environment ($\beta_{c}=1.987\pm 0.095$, $P<0.001$). These results can be interpreted as follows: when a user has a 10% higher historical preference for fast food, the odds of making a fast food order increase by $e^{\beta_{i}\times\text{0.1}}-1$=38.1%, and when the online food environment consists of 10% more fast food restaurants, the odds increase by $e^{\beta_{c}\times\text{0.1}}-1$=22.0%. Although the influence of the online food environment is relatively minor, it still has a significant association with users’ consumption behaviors, suggesting a potential negative impact of “cyber food swamps” on health outcomes.
We also observe heterogeneous effects of the online food environment on healthy food choices across different demographic groups.Figure 8 illustrates the fitted log-odds $\beta$’s for orders made by specific demographic groups. Across all groups, user preference consistently has a stronger influence than the online food environment, although the magnitude of $\beta_{c}$ varies between groups. Males, low-income individuals, and users under the age of 25 are more influenced by the environment and less by personal preference compared to their counterparts. Notably, users over the age of 40 are the least influenced by the online food environment, with $\beta_{c}=0.745\pm 0.249$, meaning that the odds of ordering fast food increase by only 7.7% when the online food environment contains 10% more fast food restaurants. This pattern aligns with the disparity trend shown in Figure 5, where groups more prone to making less healthy food delivery choices are also more affected by the online food environment. This suggests that these users may be more susceptible to online platform’s recommender system when ordering food (Sukiennik, Gao, and Li 2024; Li et al. 2022), rather than using the platform to place an order after already deciding what to eat.
@@
To identify users who changed their preferred food delivery locations during the observation period, we extract the Geohash-5 grids (approximately 5$\times$5 kilometers) in which each user was located and select those who placed more than five orders in two different grids, with each grid accounting for over 30% of their total orders. We then construct a binary time series for these users, capturing their order activity in the two grids, as shown in Figure 9. We apply a mean change detection method (Truong, Oudre, and Vayatis 2020) to identify if and when significant changes occurred in the time series, enforcing a minimum interval of 21 days between change points to avoid misdetecting fluctuations. Figure 9(a) illustrates a successful detection of a context change, where the user shifted from grid “wx552” to “wxh12” in January 2018. In contrast, the user in Figure 9(b) uniformly used O2O food delivery services across two grids, with no significant change detected. Using this method, we identify 1,892 users who experienced a significant change in their order contexts.
@@
Based on whether the user’s order context before and after the change is classified as a high fast food context (ϕ>italic-$\phi>$0.2) or a low fast food context (ϕ<italic-$\phi<$0.2), we divide users into four groups. Among them, 957 moved from a low to low fast food context, 481 from a high to high context, 208 from a low to high context, and the remaining 246 from a high to low context. To estimate the causal impact of changes in the online food environment on users’ healthy food choices, we employed a method based on Bayesian Structural Time Series (BSTS) (Brodersen et al. 2015). BSTS decomposes time series into multiple structural components to predict temporal behaviors. To assess the effect of shifting from a low to high fast food context, we use users moved from a low to low context as the control group, and those who moved from a low to high context as the experimental group. We compute the time series $\{\psi_{t}|-100<t<100\}$ for each group, where $\psi_{t}$ represents the proportion of fast food orders after $t$ days of changing context. BSTS then predicts the counterfactual outcome for the experimental group based on $\{\psi_{t}^{\text{Experiment}}|t<0\}$ by applying the temporal patterns of $\{\psi_{t}^{\text{Control}}\}$, as shown by the dashed blue lines in Figure 10. The counterfactual outcomes are then compared with the actual observations of $\{\psi_{t}^{\text{Experiment}}|t>0\}$ to derive the causal impact. A similar procedure is applied to estimate the effect of shifting from a high to low context.
@@
Figure 10: Evolution of the proportion of fast food orders (top) and cumulative differences in fast food order share (bottom) for user groups transitioning from a high fast food ratio online food environment to a low ratio context (left), and from low to high context (right). Shaded areas represent 50% confidence intervals of predicted counterfactual and cumulative effects.
@@
As shown in Figure 10, for users moving from a low to high fast food context, their preference for using O2O food delivery to order from fast food restaurants increases by 1.6%, with a cumulative fast food order share of 0.46 times higher than that of users who remained in a low fast food context. In contrast, users who moved from a high to low fast food context experienced a 7.6% decline in fast food order shar, with a cumulative effect 2.33 times lower than users who stayed in a high fast food context after 100 days. These results confirm that the online food environment affects users’ healthy food choices, and this influence is not merely a short-term response to visiting a new location but has a lasting impact. Moreover, the negative health impact of moving to high fast food context is less substantial than the positive health impact of moving to a low fast food context. This finding suggests that O2O food delivery platforms could mitigate the negative health effects of “cyber food swamps” by recommending more healthier restaurants to users, even in environments with a high concentration of fast food options.
@@ Discussion
Our findings substantiate the transformative role of O2O food delivery industry in reshaping urban residents’ healthy food choices, providing valuable insights for various stakeholders. For O2O food delivery platforms, enhancing recommendation algorithms to guide users towards healthier options is critical. This can be accomplished through personalized search features that prioritize healthier restaurants or meals, along with clear nutritional labeling, such as calorie counts or warnings about high sugar or fat content, particularly targeting demographic groups such as males, low-income individuals, and younger users with a high preference for fast food (Orfanos et al. 2009; Adams and White 2015). For restaurants, they broaden their menus to include a wider variety of healthier choices, demonstrating their commitment to both food quality and public health. For policymakers, they must address the spatial clustering of cyber food swamps with targeted regulations. For example, they could provide financial incentives and subsidies to encourage O2O food delivery restaurants that offer healthier alternatives. In addition, regulations should cap the concentration of fast food options in areas characterized by these swamps. Public health campaigns in schools and communities can significantly elevate awareness around healthy eating, a factor strongly associated with healthier dietary choices and improved health outcomes (Iqbal et al. 2021). Moreover, stricter labeling regulations, modeled after initiatives like the Affordable Care Act in the United States (Stein 2010), are essential to ensuring transparent nutritional information for O2O food delivery items.
@@
Our research also highlights the influence of online content in offline real-world social behaviors, aligning with previous studies on the web and social media (Grinberg et al. 2013; Hu, Farnham, and Talamadupula 2015; De Choudhury, Sharma, and Kiciman 2016; Carpinelli, Islind, and Óskarsdóttir 2024). Meanwhile, O2O food delivery serve as both a substitute for and an expansion of traditional offline dining environments, creating potential trade-offs and interactions that reshape conventional food accessibility (Li and Wang 2022). It remains unclear whether O2O food delivery intensifies or mitigates food swamps. In future research, we aim to compare the online and offline dining behaviors of the same group of users to better understand the unique role of online platforms in shaping dietary habits, as distinct from traditional in-person dining experiences.
@@
Our work has several limitations. First, we use whether an order is placed at a fast food restaurant as a proxy for healthy food choices. While fast food is generally associated with lower dietary quality, the nutritional content of items sold within the same restaurant can vary. Future research should assess the nutritional quality of foods purchased through O2O food delivery at the item level, such as sugar and fat intake, using more granular data. Additionally, our study focuses on pre-COVID-19 data in one large city, which may not reflect the rapid expansion of O2O food delivery coverage in the post-pandemic period (Poon and Tung 2024). Future works should extend both the temporal and spatial coverage of O2O food delivery consumption behaviors.
Ethical Statement.
All datasets collected from the OFDS platform were processed under stringent privacy protection protocols. User IDs were anonymized, and demographic attributes were injected with noise and coarsely categorized into no more than three levels for age and income. All geographic coordinates, including the locations of OFDS restaurants and delivery destinations, were recorded at the Geohash-6 level (approximately 1.2$\times$0.9 kilometers) to prevent the disclosure of individual locations. All analyses were conducted on offline storage with strict data access regulations. These precautions ensure that our analysis maintains high standards of user privacy protection.

### v3

@@ Introduction
By comparing the unhealthiness $\phi$ of offline and online food environments with fast food preferences in regional O2O food delivery orders, we find that the correlation is stronger for online environments ($r=0.204$) compared to offline environments ($r=0.092$). Figure 7 demonstrates that variations in online food environments are associated with more distinct differences in healthy food choices. In Figure 7(b), regions with lowest $\phi$ exhibit significant lower proportion of fast food orders compared with regions with $\phi>0.25$ ($P<$0.05, two-sided t-test). These findings suggest notable differences between online and traditional offline food environments. The stronger relationship between the online food environment and healthy food choices underscores the need for further investigation into their causal connections.
@@
For the online food order decision $y_{it}$ by user $i$ located in region $c_{it}$ at time $t$, we characterize the online food environment using $\phi(c_{it})$ (the proportion of accessible fast food restaurants), and user $i$’s fast food preference $\psi_{i}(t)$, as the proportion of fast food orders made by user $i$ within six months prior to time $t$. We focus on 91,879 users who have made at least 20 orders, extracting 3.86 million orders to fit a logistic regression model to estimate the effect of online food environment $\phi(c_{it})$ on the decision to choose fast food over non-fast food:
@@
where $y_{it}=1$ represents ordering fast food, $\text{logit}^{-1}(x)=\frac{e^{x}}{1+e^{x}}$ is the logistic function, $\beta_{0}$ is the fixed intercept term, and $\delta_{t}$ is a fixed effect term accounting for monthly variation. $\beta_{i}$ and $\beta_{c}$ characterize the separate effect of user preference and online food environment, respectively, as shown in Figure 8 and Figure 14.
Figure 8: Effect of user’s historical preference ($\beta_{i}$) and online food environment ($\beta_{c}$) on the probability of placing a fast food delivery order over a non-fast food delivery order for different gender, income, and age groups in 2018 (a) and 2024 (b). Error bars represent 95% confidence intervals for $\beta$’s.
The regression model reveals that both online food environment and personal preference are significantly associated with fast food choice, with personal preference exhibiting a higher log-odd coefficient ($\beta_{i}=3.227\pm 0.012$ in 2018, $\beta_{i}=3.170\pm 0.024$ in 2024, $P<0.001$, two-sided t-test) than the online food environment ($\beta_{c}=1.987\pm 0.095$ in 2018, $\beta_{c}=2.012\pm 0.342$ in 2024, $P<0.001$, two-sided t-test). These results can be interpreted as follows: when a user has a 10% higher historical preference for fast food, the odds of making a fast food order increase by $e^{\beta_{i}\times\text{0.1}}-1$=38.1% for 2018 and 37.3% for 2024, and when the online food environment consists of 10% more fast food restaurants, the odds increase by $e^{\beta_{c}\times\text{0.1}}-1$=22.0% for 2018 and 22.3% for 2024. Similar patterns are observed in Chengdu and Xiamen. Although the influence of the online food environment is relatively minor, it still has a significant association with users’ consumption behaviors, suggesting a potential negative impact of “cyber food swamps” on health outcomes.
We also observe heterogeneous effects of the online food environment on healthy food choices across different demographic groups. Figure 8 illustrates the fitted log-odds $\beta$’s for orders made by specific demographic groups. Across all groups, user preference consistently has a stronger influence than the online food environment, although the magnitude of $\beta_{c}$ varies between groups. Low-income individuals and users under the age of 25 are more influenced by the environment and less by personal preference compared to their counterparts. Notably, users over the age of 40 are the least influenced by the online food environment, with $\beta_{c}=0.745\pm 0.249$ in 2018, meaning that the odds of ordering fast food increase by only 7.7% when the online food environment contains 10% more fast food restaurants. This pattern aligns with the disparity trend shown in Figure 5, where groups more prone to making less healthy food delivery choices are also more affected by the online food environment. This suggests that these users may be more susceptible to online platform’s recommender system when ordering food (Li et al. 2022; Sukiennik, Gao, and Li 2024), rather than using the platform to place an order after already deciding what to eat.
@@
To identify users who changed their preferred food delivery locations during the observation period, we extract the Geohash-5 grids (approximately 5$\times$5 kilometers) for each order and select users who placed more than five orders across two different grids, with each grid representing more than 30% of their total orders. We then construct a binary time series for these users, capturing their order activity in the two grids, as shown in Figure 9. We apply a mean change detection method (Truong, Oudre, and Vayatis 2020) to identify if and when substantial changes occurred in the time series, enforcing a minimum spatial shift of 2 kilometers and a minimum interval of 21 days between change points to avoid mis-detecting fluctuations. Figure 9(a) illustrates a successful detection of a context change, where the user shifted from grid “wx552” to “wxh12” in January 2018. In contrast, the user in Figure 9(b) uniformly used O2O food delivery services across two grids, with no significant change detected. Using this method, we identify 1,892 users who experienced a substantial change in their order contexts in 2018. On average, the spatial shifts in their order context are 11.66 kilometers, indicating that their online food delivery environment has completely changed, as supported by Figure 1.
@@
Based on whether the user’s order context before and after the change is classified as a high fast food context (ϕ>italic-$\phi>$0.2) or a low fast food context (ϕ<italic-$\phi<$0.2), we divide users into four groups. Among them, 957 moved from a low to low fast food context, 481 from a high to high context, 208 from a low to high context, and the remaining 246 from a high to low context. To estimate the causal impact of changes in the online food environment on users’ healthy food choices, we employed a method based on Bayesian Structural Time Series (BSTS) (Brodersen et al. 2015). BSTS decomposes time series into multiple structural components to predict temporal behaviors. To assess the effect of shifting from a low to high fast food context, we take users moved from a low to low context as the control group, and those who moved from a low to high context as the experimental group. We compute the time series $\{\psi_{t}|-100<t<100\}$ for each group, where $\psi_{t}$ represents the proportion of fast food orders after $t$ days of changing context. BSTS then predicts the counterfactual outcome for the experimental group based on $\{\psi_{t}^{\text{Experiment}}|t<0\}$ by applying the temporal patterns of $\{\psi_{t}^{\text{Control}}\}$, as shown by the dashed blue lines in Figure 10. The counterfactual outcomes are then compared with the actual observations of $\{\psi_{t}^{\text{Experiment}}|t>0\}$ to derive the causal impact. A similar procedure is applied to estimate the effect of shifting from a high to low context, with users who moved from a high to high context as the control group.
@@
Figure 10: Evolution of the proportion of fast food orders (top) and cumulative differences in proportion of fast food order (bottom) for user groups transitioning from a high fast food ratio online food environment to a low ratio context (left), and from low to high context (right) in Beijing, 2018. Shaded areas represent 50% confidence intervals of predicted counterfactual and cumulative effects.
Figure 11: Evolution of the proportion of fast food orders (top) and cumulative differences in proportion of fast food order (bottom) for user groups transitioning from a high fast food ratio online food environment to a low ratio context (left), and from low to high context (right) in Beijing, 2024. Shaded areas represent 50% confidence intervals of predicted counterfactual and cumulative effects.
@@
As shown in Figure 10, for users moving from a low to high fast food context, their preference for using O2O food delivery to order from fast food restaurants increases by 1.6%, with a cumulative proportion of fast food order of 0.46 times higher than that of users who remained in a low fast food context. In contrast, users who moved from a high to low fast food context experience a 7.6% decline in their proportion of fast food orders, with a cumulative effect 2.33 times lower than users who stayed in a high fast food context after 100 days. Similar trends were observed in 502 users who shifted their online food environment in 2024, as depicted in Figure 11. Specifically, moving from a low to high fast food context leads to a 0.2% increase in the proportion of fast food orders, while shifting from a high to low fast food context results in a 9.6% decrease in the proportion of fast food orders. These results confirm that the online food environment affects users’ healthy food choices, and this influence is not merely a short-term response to visiting a new location but has a lasting impact. Moreover, the negative health impact of moving to high fast food context is less substantial than the positive health impact of moving to a low fast food context. This finding suggests that O2O food delivery platforms could mitigate the negative health effects of “cyber food swamps” by recommending much healthier restaurants to users, even in environments with a high concentration of fast food options.
@@ Discussion
Our findings substantiate the transformative role of O2O food delivery industry in reshaping urban residents’ healthy food choices, providing valuable insights for various stakeholders. For O2O food delivery platforms, enhancing recommendation algorithms to guide users towards healthier options is critical. This can be accomplished through personalized search features that prioritize healthier restaurants or meals (Vanderlee and Sacks 2023), along with clear nutritional labeling, such as calorie counts or warnings about high sugar or fat content (Jiang et al. 2019; Greenthal et al. 2023), particularly targeting demographic groups such as males, low-income individuals, and younger users with a high preference for fast food (Orfanos et al. 2009; Adams and White 2015). Restaurants stakeholders can broaden their menus to include a wider variety of healthier choices, demonstrating their commitment to both food quality and public health (Vanderlee and Sacks 2023). Policy makers must address the spatial clustering of cyber food swamps with targeted regulations. For example, they could provide financial incentives and subsidies to encourage O2O food delivery restaurants that offer healthier alternatives (Shill et al. 2012). In addition, regulations should cap the concentration of fast food options in areas characterized by these swamps. Public health campaigns in schools and communities can significantly elevate awareness around healthy eating, a factor strongly associated with healthier dietary choices and improved health outcomes (Iqbal et al. 2021). Moreover, stricter labeling regulations, modeled after initiatives like the Affordable Care Act in the United States (Stein 2010), are essential to ensuring transparent nutritional information for O2O food delivery items.
@@
Our research also highlights the influence of online content in offline real-world social behaviors, aligning with previous studies on the web and social media (Grinberg et al. 2021; Hu, Farnham, and Talamadupula 2021; De Choudhury, Sharma, and Kiciman 2016; Carpinelli, Islind, and Óskarsdóttir 2024). Meanwhile, O2O food delivery serve as both a substitute for and an expansion of traditional offline dining environments, creating potential trade-offs and interactions that reshape conventional food accessibility (Li and Wang 2022). It remains unclear whether O2O food delivery intensifies or mitigates food swamps. In future research, we aim to compare the offline and online dining behaviors of the same group of users to better understand the unique role of online platforms in shaping dietary habits, as distinct from traditional in-person dining experiences.
@@
Our work has several limitations. First, we use whether an order is placed at a fast food restaurant as a proxy for healthy food choices. While fast food is generally associated with lower dietary quality, the nutritional content of items sold within the same restaurant can vary. Future research should assess the nutritional quality of foods purchased through O2O food delivery at the item level, such as sugar and fat intake, using more granular data. Additionally, the uncertainty in user demographic profiles could impact the validity of the observed differences across demographic groups. However, the platform reports that the inferred demographic profiles have an accuracy of over 90%, and we filter out users with low confidence levels to ensure the reliability of our findings. Furthermore, other observations that do not involve comparisons across demographic groups remain robust and unaffected by this potential uncertainty.
Acknowledgements
This work is supported by the National Key Research and Development Program of China under grant No. 2024YFC3307600. This work is supported by the National Natural Science Foundation of China under grant No. 23IAA02114 and 62472241. This work is supported in part by Tsinghua University-Toyota Research Center.
Appendix A Appendix
Figure 12: Scaling laws of O2O food delivery restaurants (a), fast food restaurants (b), O2O food delivery orders (c), and fast food orders (d) in Chinese cities in 2024.
Figure 13: The average fast food order ratio and order price across different gender, income level, and age groups in Beijing (a), Chengdu (b), and Xiamen (c) in 2024. Error bars represent 95% confidence intervals.
Figure 14: Effect of user’s historical preference ($\beta_{i}$) and online food environment ($\beta_{c}$) on the probability of placing a fast food delivery order over a non-fast food delivery order for different gender, income, and age groups in Chengdu (a) and Xiamen (b). Error bars represent 95% confidence intervals for $\beta$’s.